a large multiplex movie house has many theaters. the largest has 35 rows. there are 20 seats in the first row. each row has two seats more than the previous row. how many total seats are there in the theater??
thanks for the help

Answer by stanbon(75887)

a large multiplex movie house has many theaters. the largest has 35 rows. there are 20 seats in the first row. each row has two seats more than the previous row. how many total seats are there in the theater??
---------------------
Find the sum of this arithmetic progression.
a(1) = 20
common difference = 2
a(35) = a(1)+ (35-1)d = 20 + 34*2 = 20+78 = 98
------
Ans: S(35) = (n/2)(a(1)+a(35)) = (35/2)(20+98) = 35*59 = 1365
----------------
Cheers,
Stan H.
-------------